OpenStudy (anonymous):

his graph of a function is a translation of y = 4/x . What is an equation for the function?

OpenStudy (anonymous):

is this a completer question???

OpenStudy (mertsj):

It would be most helpful if you would post the entire question.

OpenStudy (mertsj):

yes. I'm still trying to figure out what the question is.

OpenStudy (mertsj):

Consider the hyperbola y = 4/x If you change it to y = 4/(x+1) that moves the center 1 space to the left. If you change it to y = 4 (x-1) that moves the center 1 space to the right.

OpenStudy (mertsj):

If you change it to: \[y=\frac{4}{x} +1 \] that moves the center one space up.

OpenStudy (mertsj):

If you change it to \[y=\frac{4}{x}-1\] that moves the center one space down

OpenStudy (mertsj):

So you need to figure out where the center of the hyperbola is and then you will know how it has been moved from (0,))

OpenStudy (anonymous):

i really dnt know non of thsi stuff do u know which one of the answers it would be? @Mertsj

OpenStudy (mertsj):

The center of the hyperbola shown in the picture is (-6,2) Choose the equation with that center. And you really do need to learn this stuff.

OpenStudy (anonymous):

ok. I think its B. @Mertsj

OpenStudy (mertsj):

Choice b has been moved left 2 and down 6. so do you think that's what happened in the picture?

OpenStudy (anonymous):

no I think left 6 and down 2

OpenStudy (anonymous):

So D?

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

How is it down 2?

OpenStudy (anonymous):

left 6 up 2

OpenStudy (anonymous):

now im confused so its not B or D so last options are A and C... I don't think its A

OpenStudy (mertsj):

Yes. Left 6 and up 2 so which equation is that?

OpenStudy (mertsj):

Well then, if it is not A, B or D which one is it?

OpenStudy (anonymous):

C lol

OpenStudy (anonymous):

But is that correct?

OpenStudy (mertsj):

Yes.

OpenStudy (anonymous):

awesome thank you!!! <3

OpenStudy (mertsj):

yw